-2(3x-1)=4(3x-2)-x

Found 2 solutions by xkatiexprime, rfer:
Answer by xkatiexprime(4)

ever heard of distrubuting? WELL. let's start with the left side of the equal sign...
-2(3x - 1)
you multiply the -2 by all the numbers and stuff in the brackets
= -6x + 2
now with the right side of the = sign
4(3x - 2) - x
distribute the 4 into the numbers in the bracket
= 12x - 8 - x
therefore the new equation is -6x + 2 = 12x - 8 - x
separate the variable collect the like terms
2 + 8 = 12x + 6x - x
10 = 17x
0.59 = x
you can always check your answer by subbing your answer into the equation
-2(3x -1) = 4(3x - 2) - x
-2[3(0.59) - 1] = 4[3(0.59) - 2] - 0.59
-2[1.77 - 1] = 4[1.77 - 2] - 0.59
-2[0.77] = 4[-0.23] - 0.59
-1.54 = -0.92 - 0.59
-1.54 = -1.54
yay! they match! x is about 0.59. :)

Answer by rfer(16322)

-6x+2=12x-8-x
-6x-11x=-8-2
-17x=-10
x=-10/-17
x=10/17
x=0.588